# Language Model Council: Democratically Benchmarking Foundation Models on Highly Subjective Tasks

### v2

@@ C.1 Invariability
C.2 Pairwise Positional Consistency (PPC)
@@ C.3 Experiment
For a single pair of responses, there are 10 reps (5 reps for each prompt * 2 prompts) in one order and 10 reps in the swapped order. Thus, there are $10*10=100$ possible rating couplets, which forms the denominator for PPC calculation.
@@
    Coarse preferences with tie option (A>B, B>A, Asimilar-$\sim$=B)
@@
    Coarse preferences without tie option (A>B, B>A)
  3. 3.
    Granular preferences with tie option (A>>B, A>B, B>A, B>> A, Asimilar-$\sim$=B)
  4. 4.
    Granular preferences without tie option (A>>B, A>B, B>A, B>>A)
C.4 Results
temperature=0 is the best for reliability. The best average invariability and pairwise positional consistency across all 20 LLMs on the council is achieved with $temperature=0$. To our surprise, only $13/20$ models produce fully invariant ratings, even with $temperature=0$.
Coarse or granular rating options? Under $temperature=0$, the invariability between using coarse and granular rating options is small (0.98 vs. 0.96). The difference in PPC is more stark (0.91 vs. 0.80), though still tolerable. We decide to proceed with granular rating options for the main experiment for parity with Arena Hard Li et al. (2024a) to give more weight to strong preferences in the ELO calculation.
Figure 16 shows calibration scores for invariability and PPC, averaged over 20 LLMs and 10 repetitions for each under different pairwise comparison options. Table 11 shows a detailed breakdown per LLM, using granular pairwise comparison options without ties and with $temperature=0$.
C.5 Conclusion
@@
Appendix D Details on Reference Model Selection for the EI Case Study
@@
D.1 Understanding the Bradley-Terry Procedure
@@
D.2 What is the reference model?
@@
D.3 Selecting our reference model
In our arena-based LMC EI case study, we use a single reference model, following the approach of other arena-based benchmarks like Chatbot Arena Hard Li et al. (2024a) and Alpaca Eval Dubois et al. (2024a), which used GPT-4-0314 and GPT-4-turbo, respectively. It is unclear how Li et al. (2024a) and Dubois et al. (2024a) chose their reference models, but we will explain our choice.
@@ D.3 Selecting our reference model
Appendix E Human Evaluation
@@
All participants are over 18 years old. Our sample is made up of 53 women, 46 men, and one non-binary identifying individual. 84 of our participants were from the United Kingdom, 14 from the United States and two from other English-speaking countries; all were native English speakers. With regards to their use of AI chatbots, 23 report using them every day or nearly every day, 48 sometimes, four rarely and only four report never using them. None report having difficulties reading long texts. None report having difficulties reading long texts.
@@
Appendix F More Details on Comparison to Other Leaderboards
Figure 23: Extension of Figure 4 with more Spearman correlations with our EI Human Study (Appendix E) for benchmark scores for 9 LLMs (listed in Figure 3). We include the correlation scores for all individual LLM judges on the LMC, as well as correlations for hand-curated sub-councils discussed in Section 5.2.
F.1 Mining an EQ subset of Chatbot Arena
Chatbot Arena includes many prompts that may have little to do with EI, so it may be unsurprising that the overall correlation with our human study is low (0.48) (Figure 4).
In this section, we outline a procedure to produce an EI-based re-ranking of LLMs based on an EI-targeted subset of Chatbot Arena prompts.
A publicly released subset of Chatbot Arena’s prompts are available online [^15] https://huggingface.co/datasets/lmsys/chatbot_arena_conversations. This dataset has 33K unique prompts. Because there are no fine-grained categories that allow us to easily slice the leaderboard by performance on specific questions, we send all 33K unique prompts to a reasonably competent LLM, llama-3.1-8b Meta (2024), one prompt at a time, to assess whether the prompt is EI-related or not. The prompt template used to perform a classification of EI-relatedness is in Figure 24.
@@ F.1 Mining an EQ subset of Chatbot Arena
This results in 2680 prompts ( 8%) flagged to be potentially useful for assessing EI. While a stronger LLM could be used to do this EI-relatedness flagging, most of the examples we spot-checked looked reasonably related to EI. Here are some examples:
@@
  • “Hey I’m a little stressed. I need some easy ways to relax”
@@
Following the same procedure as the LMC, we assess responses in a pairwise fashion with position flipping using Qwen-1.5-32B’s responses as the reference and with GPT-4o-mini OpenAI (2024b) as the judge, mimicking the single judge design of Chatbot Arena Hard Li et al. (2024a). This results in 100 x 8 x 2 = 1600 ratings.
@@
Appendix G Qualitative Analysis: What Makes a Response Preferred Over Another?
G.1 Motivation
Several arena-based benchmarks (ours included) have demonstrated that a clear ranking among LLMs can be established, but there is not much understanding as to why the rankings are the way they are. For example, platforms like Chatbot Arena do not clarify how factors like feel and style are weighed against correctness Wei (2024), and while many evaluation systems like AlpacaEval Dubois et al. (2024a) or MT-Bench Zheng et al. (2024) tout chain-of-thought (CoT) prompting Wei et al. (2022) to improve the explainability of ratings by LLM judges, these justifications are left unanalyzed.
G.2 Reasoning trace themes extraction procedure
G.3 Subjectivity of defining and assigning themes
  1. 1.
  2. 2.
  3. 3.
G.4 Results and discussion
@@
more suggestions … less verbose -0.144
@@ G.5 Conclusion
Appendix H Quantifying the Stability of a Benchmark
@@
H.1 Mathematical Definition of Mean Expected Rank Variance (MERV)
@@ H.2 Understanding MERV
H.3 Comparison with Separability
Separability, in contrast, measures the percentage of respondent pairs with completely non-overlapping confidence intervals, often derived from bootstrapping Li et al. (2024a). It quantifies the statistical significance of performance differences, focusing on how distinct the rankings of respondents are in terms of performance intervals.
@@ H.3 Comparison with Separability
Appendix I Prompt Templates
@@
Figure 31: Prompt variations on Figure 30 (applied to the bottom highlighted text) used to study natural consistency and variability under different pairwise comparison regimes in Appendix C.

### v4

@@ B.1 Invariability
B.2 Pairwise Positional Consistency (PPC)
@@ B.3 Experiment
For a single pair of responses, there are 10 repetitions (5 repetitions for each prompt * 2 prompts) in one order and 10 reps in the swapped order. Thus, there are $10*10=100$ possible rating couplets, which forms the denominator for the calculation of PPC.
@@
  • $\bullet$
    Coarse preferences with tie option
    (A>B, B>A, A=B)
  • $\bullet$
    Coarse preferences without tie option
    (A>B, B>A)
  • $\bullet$
    Granular preferences with tie option
    (A>>B, A>B, B>A, B>> A, A=B)
  • $\bullet$
    Granular preferences without tie option
    (A>>B, A>B, B>A, B>>A)
B.4 Results
temperature=0 is superior for reliable and consistent judgments. The best average invariability and pairwise positional consistency across all 20 LLMs on the council is achieved with $temperature=0$. To our surprise, only $13/20$ models produce perfectly invariant ratings across all repetitions, even with $temperature=0$.
Coarse or granular rating options? Under $temperature=0$, the difference in invariability between using coarse and granular rating options is small (0.98 vs. 0.96). The difference in PPC is more stark (0.91 vs. 0.80), though still tolerable. We decide to proceed with granular rating options for the main experiment to maintain parity with Arena Hard Li et al. (2024b) and to give more weight to strong preferences in the final ELO calculation.
Full findings. Figure 16 shows calibration scores for invariability and PPC, averaged over 20 LLMs and 10 repetitions for each under different pairwise comparison options. Table 11 shows a detailed breakdown per LLM, using granular pairwise comparison options without ties and with $temperature=0$.
B.5 To CoT or not to CoT?
Newer research suggests that Chain-of-Thought (CoT) prompting may degrade LLM performance on non-math and non-symbolic reasoning tasks, which may include ratings for simple pairwise comparisons Sprague et al. (2024).
We assert our use of CoT prompting for two main reasons:
@@ B.5 To CoT or not to CoT?
    It aligns with conventions in prior literature and arena-based LLM evaluation settings Li et al. (2024b); Chiang et al. (2024)..
@@
    CoT prompts generate reasoning traces, which we analyzed (Appendix F) to better understand the rationale behind the rankings.
B.6 Conclusion
@@
Appendix C Details on Reference Model Selection for the EI Case Study
@@
C.1 Understanding the Bradley-Terry Procedure
@@
C.2 What is the reference model?
@@
C.3 Selecting our reference model
In our arena-based LMC EI case study, we use a single reference model, following the approach of other arena-based benchmarks like Chatbot Arena Hard Li et al. (2024b) and Alpaca Eval Dubois et al. (2024), which used GPT-4-0314 and GPT-4-turbo, respectively. It is unclear how Li et al. (2024b) and Dubois et al. (2024) chose their reference models, but we will explain our choice.
@@ C.3 Selecting our reference model
Appendix D Human Evaluation
@@
All participants are over 18 years old. Our sample is made up of 53 women, 46 men, and one non-binary identifying individual. 84 of our participants were from the United Kingdom, 14 from the United States and two from other English-speaking countries; all were native English speakers. With regards to their use of AI chatbots, 23 report using them every day or nearly every day, 48 sometimes, four rarely and only four report never using them. None report having difficulties reading long texts.
@@
Appendix E More Details on Comparison to Other Leaderboards
Figure 23: Spearman ranking correlations with our EI Human Study (Appendix D) for benchmark scores for 9 LLMs (listed in Figure 3). We include the correlation scores for all individual LLM judges on the LMC, as well as correlations for hand-curated sub-councils discussed in Section 5.2.
Figure 24: Kendall-Tau ranking correlations with our EI Human Study (Appendix D) for benchmark scores for 9 LLMs (listed in Figure 3). We include the correlation scores for all individual LLM judges on the LMC, as well as correlations for hand-curated sub-councils discussed in Section 5.2.
Mining an EQ subset of Chatbot Arena
Chatbot Arena includes many prompts that may have little to do with EI, so it may be unsurprising that the overall correlation with our human study is low (0.48) (Figure 4). In this section, we outline a procedure to produce an EI-based re-ranking of LLMs based on an EI-targeted subset of Chatbot Arena prompts.
A publicly released subset of Chatbot Arena’s prompts are available online. [^15] https://huggingface.co/datasets/lmsys/chatbot_arena_conversations This dataset has 33K unique prompts. Because there are no fine-grained categories that allow us to easily slice the leaderboard by performance on specific questions, we send all 33K unique prompts to a reasonably competent LLM, llama-3.1-8b Meta (2024), one prompt at a time, to assess whether the prompt is EI-related or not. The prompt template used to perform a classification of EI-relatedness is in Figure 25.
@@ Appendix E More Details on Comparison to Other Leaderboards
This results in 2680 prompts (~8%) flagged to be potentially useful for assessing EI. While a stronger LLM could be used to do this EI-relatedness flagging, most of the examples we spot-checked looked reasonably related to EI. Here are some examples:
  • $\bullet$
@@
Following the same procedure as the LMC, we assess responses in a pairwise fashion with position flipping using Qwen-1.5-32B’s responses as the reference and with GPT-4o-mini OpenAI (2024b) as the judge, mimicking the single judge design of Chatbot Arena Hard Li et al. (2024b). This results in 100 x 8 x 2 = 1600 ratings.
@@
Appendix F Qualitative Analysis: What Makes a Response Preferred Over Another?
@@
more suggestions less verbose -0.144
@@
F.1 Motivation
Several arena-based benchmarks (ours included) have demonstrated that a clear ranking among LLMs can be established, but there is not much understanding as to why the rankings are the way they are. For example, platforms like Chatbot Arena do not clarify how factors like feel and style are weighed against correctness Wei (2024), and while many evaluation systems like AlpacaEval Dubois et al. (2024) or MT-Bench Zheng et al. (2023) tout chain-of-thought (CoT) prompting Wei et al. (2022) to improve the explainability of ratings by LLM judges, these justifications are left unanalyzed.
F.2 Reasoning trace themes extraction procedure
F.3 Subjectivity of defining and assigning themes
  • 1.
  • 2.
  • 3.
F.4 Results and discussion
@@ F.5 Conclusion
Appendix G Quantifying the Stability of a Benchmark
@@
G.1 Mathematical Definition of Mean Expected Rank Variance (MERV)
@@ G.2 Understanding MERV
G.3 Comparison with Separability
Separability, in contrast, measures the percentage of respondent pairs with completely non-overlapping confidence intervals, often derived from bootstrapping Li et al. (2024b). It quantifies the statistical significance of performance differences, focusing on how distinct the rankings of respondents are in terms of performance intervals.
@@ G.3 Comparison with Separability
Appendix H Prompt Templates
@@
Figure 32: Prompt variations on Figure 31 (applied to the bottom highlighted text) used to study natural consistency and variability under different pairwise comparison regimes in Appendix B.